# Phase segregation lubricates paste extrusion

### v1

@@ III The tribological role of the lubrication layer during pellet extrusion
Figure 3: Lubrication layer affects the extrusion energy balance. a, Lubrication reduces the apparent die wall friction $\mu$ and, therefore, (b) reduces the net specific mechanical energy $(\mathrm{SME_{net}})$ use with increasing lubrication layer thickness $\lambda$. The reduction in mechanical energy use in particular (c) reduces die heating through friction and consequently affects the change in pellet temperature $\Delta T_{\rm die}$ over the die—the difference between the pellet’s temperature after extrusion and the ingredients temperature before extrusion, as well as final pellet temperature. The horizontal error-bars in (b) and (c) correspond to the minimum and maximum values for $\lambda$ as represented in Fig. 1d. The vertical error bars represent minimum-maximum range of at least three samples during experimental runs (see Methods section).
IV Conclusions
@@
We confirmed the existence and thickness of a micrometer-thick lubrication layer that forms via particle-fluid phase segregation during the high pressure extrusion of pastes. We show that this lubrication layer directly affects the tribological properties of the paste flow. We used realistic, plant-based materials inside a an industrial-scale pellet extruder to evidence how the lubrication layer interlinks process variables during industrial extrusion. We so unravel the fundamental mechanisms behind a decades old observation on the lubricating effect of water addition during pellet production [37]. Our work reveals that the friction within the die is directly influenced by the thickness of the lubrication layer, providing insights for future monitoring of product quality, shape, and performance in food, feed, and pharmaceutical extrusion processes and adjacent extrusion disciplines which process inorganic ingredient streams like ceramics. Our work provides a new framework to understand and use the role of other process parameters such as die geometry, particle size distribution, and ingredient composition. Moreover, continuous monitoring of the extrusion process offers a novel approach for process control, with immediate practical implications for optimizing industrial production on a global scale. Our findings can lead to significant energy reductions, contribute to a more sustainable extrusion process and enhanced quality control in the production of a wide range of extrusion-based materials.
@@ Acknowledgements.
We gratefully acknowledge discussions with the “Pelleting in Circular Agriculture” consortium members. We thank the Wageningen Electron Microscopy Center (WEMC) of Wageningen University & Research for access to facilities We thank the Van der Waals-Zeeman Institute for providing access to the Keyence VK-X1000. We thank Zorheh Farmani and Jordi Rijpert for their help during the experimental investigation. We thank Remco Fokkink for helping with the development of software to read out the relevant sensors during the drying experiments. The authors thank Synchrotron SOLEIL for the ANATOMIX beamline time (proposal n° 20220272). ANATOMIX is an Equipment of Excellence (EQUIPEX) funded by the Investments for the Future program of the French National Research Agency (ANR), project NanoimagesX, grant no. ANR-11-EQPX-0031.
@@
  • Funding: this study was financially supported by The VICTAM Foundation, Agrifirm NWE B.V., DSM, Elanco Animal Health, Pelleting Technology Netherlands, Phileo S.I. Lesaffre, Topsector Agri & Food, Zetadec, and Wageningen University & Research, as partners in the project “Pelleting in Circular Agriculture” (project number: LWV1965)
@@ Appendix A Extended Data
Figure A.1: Drying experiments differentiate between fresh (quick drying) and rehydrated (slow drying) pellets, complemented by numerical results. The effect of radial phase segregation on pellet drying behavior was evaluated both experimentally and numerically. We compared the drying behavior of rehydrated pellets with a uniform moisture distribution and freshly produced pellets with increased surface wetness. a, A schematic representation illustrates the initial radial moisture distribution used as the initial condition in numerical modeling of the drying process for freshly produced pellets. The moisture ratio (unitless) represents the normalized mass decay over $900$ seconds of drying (see Eq. 6). Numerical simulations provide a qualitative comparison of the effect of radial phase segregation on drying behavior, shown in both numerical simulations (b) and experimental drying data (c) for fresh (blue lines) and rehydrated pellets (yellow lines) using the moisture distributions as the initial condition (a). The difference in drying behavior is caused by radial phase segregation, specifically the formation of a lubrication layer during pellet extrusion.
We examined the pellet drying behavior, internal and external pellet microstructure, and surface topology of pellets produced under systematically varied process conditions, adjusting only the amount of added steam to the ingredient mixture when formulating the paste, thereby changing its initial water content prior to compaction, $\Delta\mathrm{mc}$. Our objective was to determine whether phase segregation occurs during pellet extrusion. In the processing of pastes and other dense particulate suspensions, such phase segregation processes are also referred to as “spatial demixing,” “liquid migration,” “phase redistribution,” “drainage,” or “self-filtration” [11, 33]. According to our hypothesis, a thin lubricating film forms between the pellet and the metal die due to phase segregation (Fig. 1a), which fills the asperities between the two sliding surfaces and reducing the frictional contact area. As a result of the radial phase migration, the moisture distribution across the pellets becomes non-uniform, with the exterior surface of the pellet facing the die wall experiencing increased moisture levels during extrusion compared to the pellet’s interior volume [22]. Through the combined circumstantial evidence presented below, we conclude that such phase migration occurs during pellet extrusion.
@@ A.1 Pellet production
A.2 Water distribution inside pellets
The concept of phase segregation during extrusion of multi-phase ingredient mixtures is not unique to pellet extrusion and occurs under a wide range of process conditions. Ranging from low viscosity soft matter suspensions, e.g. in food [50] or human health [11] applications, as well as during the extrusion of highly viscous flows, like rock melts [51]. As a result of the physical forces, an uneven distribution of the liquid fraction occurs across the product during or after extrusion. However, direct analysis of liquid distribution inside extrusion channels using non-invasive techniques like NMR and MRI at industrial process conditions is currently not possible.
Extrusion processes occur within metal dies typically made of hardened stainless steel, which can withstand high processing temperatures and pressures. This prevents the use of non-invasive experimental techniques [29, 31], along with the significant amount of dust produced in pelleting facilities which interferes with the non-invasive experimental techniques.
To address these limitations, we conducted a numerical evaluation of how radial redistribution of water over the pellet influences drying characteristics immediately after extrusion. The numerical results were qualitatively compared with experimental drying data obtained in the pilot plant.
In the numerical simulations, the initial moisture distribution along the radial coordinate $r$ was governed by a half-Gaussian function $M(r)$—a Gaussian centered at the pellet-air interface ($r_{p}$).
$M(r)=M_{\rm surf}\exp\left[-\frac{1}{2}\left(\frac{r-r_{p}}{\sigma}\right)^{2}\right]$ (1)
@@ A.2 Water distribution inside pellets
Consequently, for the numerical model, mass transfer coefficient ($k_{c}$) and effective diffusion coefficient ($D_{\rm eff}$) were selected in a way that the transport rate at which water moves within the pellet (diffusion of water) is much lower than the rate of evaporation (mass transfer rate representing the change of liquid water on the surface into water vapor within the passing air). This is associated with a mass Biot number ($\mathrm{Bi_{m}}$) much larger than $1$ (see Appendix C.1). With a mass Biot number much larger than $1$, we ensure that the rate of evaporation at the surface does not limit the drying process, as the maximum evaporation rate is lower than the vapor capacity of the passing fresh air, as observed in the experimental data (see Fig. 1b).
@@
The qualitative comparison between the numerical model and the experimental drying data supports the hypothesis that phase segregation occurs during pellets extrusion. Combining the quick initial drying of the fresh pellets together with the observation of air saturation (Fig. 1 b), we conclude that water must be present in close proximity to the pellet exterior surface as a result of phase segregation during pellet extrusion.
@@ A.3 Pellet microstructure
During pellet extrusion, the finely ground organic ingredients are bound into a cylindrical shaped pellet. Such organic ingredient streams deform easily at the typical process pressures, resulting in densification of the ingredient mixture and subsequently enhances the agglomeration process through various bond types [20]. The mechanical properties [57, 58], as well as the glass and melt transition temperatures of such organic plant derived products are strongly influenced by the local moisture content [6, 36]. Consequently, the local increase in moisture content due to the presence of a water rich lubrication layer at the pellets surface is expected to affect the surface microstructure of the pellet. Hence, we employed x-ray tomography at the synchrotron SOLEIL, utilizing the ANATOMIX beamline [59], to analyze the pellet microstructure. This analysis focused on identifying density gradients and structural changes associated with shear deformations and moisture gradients. Such gradients and structural changes can offer insights into the flow profile inside the extrusion die and the local moisture content during extrusion.
@@ A.3.1 A uniform radial pellet density profile
Figure A.2: The pellets exhibit uniform radial density. Radial density profiles of pellets with a diameter of 6 mm, obtained through analysis of x-ray tomograms obtained at the ANATOMIX beamline, for two process conditions with (a) $\Delta\mathrm{mc}$ $0.009$ (low water content) and (b) $0.057$ (high water content) respectively. The light gray lines in (a) and (b) correspond to individual slices along the 3D volume. The black lines in (a) and (b) correspond to the average density profile over each of the slices within the stack (see Materials & Methods). The region marked by the gray area, corresponds to the area in which we observe a phase-contrast enhancement artifact. In this region the apparent density could not be determined accurately.
@@ A.3.3 Internal pellet deformation
It has been previously proposed that the ingredient stream moves as a plug inside the die channels [60], with a uniform flow velocity through the pellet die. A lubrication layer should facilitate such plug-flow and therefore, signs of internal deformations caused by velocity gradients should be absent. A qualitative analysis of x-ray tomograms confirms this absence of internal deformations along the longitudinal axis as indications of the flow direction (Fig. A.5). Additionally, the overall porosity is low (Fig. A.5), suggesting that the ingredients are forced into place, approaching unit density during pelletization. Indeed, the estimated process pressures during pelletization [61], are of similar order, or larger than the typical compressive strength of the particles [62] and hence the particles are compressed into a tightly packed pellet with a low porosity. These observations support the hypothesis of phase segregation, wherein water in between the particles, is expelled from the ingredient matrix into the pellet-die gap during pellet extrusion, leading to the formation of a lubrication layer that enables plug flow inside the pellet die.
Figure A.5: Longitudinal cross sections (x-ray tomograms obtained at the ANATOMIX beamline) of four pellets produced during the experimental trial at two different water contents after steam treatment, (a) two at low $\Delta\mathrm{mc}$ and (b) two at high $\Delta\mathrm{mc}$, 0.009 and 0.057 kg water per kg ingredients respectively. No indication for the flow direction is visible within the cross sections. The pellets flow as plugs through the pelleting die [60]. Scale bar indicates $0.5$ cm.

### v2

@@ Introduction
Figure 4: (a) Lubrication reduces the apparent die wall friction $\mu$ and, therefore, (b) reduces the net specific mechanical energy $(\mathrm{SME_{net}})$ use with increasing lubrication layer thickness $\lambda$. The reduction in mechanical energy use in particular (c) reduces die heating through friction and consequently affects the change in pellet temperature $\Delta T_{\rm die}$ over the die—the difference between the pellet’s temperature after extrusion and the ingredients temperature before extrusion, as well as final pellet temperature. The horizontal error-bars in (b) and (c) correspond to the minimum and maximum values for $\lambda$ as represented in Fig. 2d. The vertical error bars represent minimum-maximum range of at least three samples during experimental runs (see Appendix B).
Friction dominates extrudate flow dissipation
So far, we have only provided evidence of the existence of a liquid surface layer on the extrudate. We can confirm the lubricating effect of this layer by first establishing that energy use in the extrusion of pastes is set predominantly by friction of the extrudate inside the die channel. As pelleting trial do not allow us to change the applied pressure that drives extrusion, we verify the importance of friction in extrusion by probing how mechanical energy dissipation depends on the die geometry. Relying on the same physical principles as the Hagen-Janssen effect [48], Holm et al. [49, 50], derived a relationship that couples compression ratio - die length over die diameter, $L/D$ (-), and the coefficient of friction, $\mathrm{\mu}$, to the pressure required for extrusion during pelletization:
$P_{x}(x)=\frac{P_{0}}{\nu}\exp\left(4\mu\nu\frac{L}{D}-1\right)$ (1)
In which the extrusion pressure, $P_{x}$ $(\mathrm{Pa})$, is determined by a prestressing constant, $P_{0}$ $(\mathrm{Pa})$, the radial-to-axial stress ratio constant, or lateral pressure coefficient, $\nu$ (-), the coefficient of friction and the compression ratio of the die (see [49, 51] for further reference). By taking $\nu$ as an ingredient specific constant during pelletization, the pressure required for extrusion increases exponentially with $\mu$ and $L/D$. Furthermore, the extrusion pressure $(\mathrm{Nm^{-2}})$ and power consumption $(\mathrm{N\ ms^{-1}=W})$ are related through the volume flow rate $(\mathrm{m^{3}s^{-1}})$. Consequently, the power consumption of the pellet press, as well as the specific mechanical energy usage - the energy per unit mass - is expected to increase exponentially as a function of $\mu$ and $L/D$ as well, which we confirm in Fig. 3.
[…]
Energy reduction due to lubrication
As we have seen, the sliding friction between the pellet and the metal die contributes to the energy use of the pellet press. This energy use is proportional to the extrusion pressure and increases exponentially with $L/D$ and with the apparent coefficient of friction between the die and the paste. The apparent coefficient depends on the solid friction coefficient and is determined by the paste composition [52, 53]. This dry coefficient of friction can not be actively manipulated during pellet extrusion, yet the formation of a thin lubrication layer can contribute to a smooth transition from the full contact “boundary lubrication regime” into the so-called “mixed regime” [28], in which part of the solid contact is replaced by a lubrication layer, significantly reducing the coefficient of friction (see appendix A.4).
The decrease in the apparent coefficient of friction, $\mu$, at the die-wall surface has a direct impact on the specific mechanical energy (SME) use of the pellet press—the amount of mechanical energy required to produce one kilogram of pellets. This is exemplified in Fig. 4b, which shows that increasing the lubrication layer thickness, due to an increase in the initial $\phi_{\rm liquid}/\phi_{\rm solid}$ ratio after steam conditioning, reduces the specific mechanical energy (SME) during extrusion by approximately 50%. Hence, adjusting the lubrication layer through steam addition enables modification of the friction coefficient within a running production line, without altering the ingredient mixture composition or press/die configuration. In addition to the change of the initial $\phi_{\rm liquid}/\phi_{\rm solid}$ ratio, the extrusion rate is another key factor that modulates friction and energy dissipation during pellet extrusion. An increase in extrusion rate reduces frictional resistance and thus lowers SME, in accordance with Stribeck curve expectations [28] (see appendix Fig. A.7).
Overall, these changes at the die-wall boundary affect the amount of heat dissipated during extrusion. An increase in lubrication layer thickness reduces the heating of the pellet during extrusion, here quantified by $\Delta T_{\rm die}$. This effect is in Fig. 4c, where the increasing in lubrication layer thickness is linked to a reduction in $\Delta T_{\rm die}$ due to the reduced die-wall friction. These changes in the extrudate’s final temperature also affect final product properties, such as “durability”, as detailed in appendix A.5, and elsewhere [23, 22]. We reason that these temperature fluctuations change the physico-chemical properties of the organic ingredient streams, which may result in gelatinization and melting of starches and denaturation of proteins, which subsequently affect the nutritional quality and availability during consumption [39]. Measurement and subsequently tuning of the lubrication layer is thus essential for further optimization of pelleting processes as it enables a more precise control over the mechanical energy use and product temperature compared to the current set of operating variables.
[…]
Numerical evidence for $\lambda$
The concept of liquid migration during extrusion of multi-phase ingredient mixtures is not unique to extrusion and occurs under a wide range of process conditions. Ranging from low viscosity soft matter suspensions, e.g. in food [54] or human health [12] applications, as well as during the extrusion of highly viscous flows, like rock melts [55]. As a result of the physical forces, an uneven distribution of the liquid fraction occurs across the product during or after extrusion. However, direct analysis of liquid distribution inside extrusion channels is currently not possible. The higher pressure extrusion conditions make non-invasive techniques like Magnetic Resonance Imaging or Optical Coherence Tomography incompatible or do not have the required spatial resolution [32, 34].
To address these limitations, we also conducted a numerical evaluation of how radial redistribution of water over the pellet influences drying characteristics immediately after extrusion. The numerical results were qualitatively compared with experimental drying data obtained in the experiment. We describe the numerical setup at length in Appendix A. Our modeling demonstrates that the saturation of the drying air and the initial rate of drying, observed during the drying of freshly produced pellets, can only occur if the evaporating water comes directly from the pellet surface. This is in contrast to pellets with an elevated but uniform moisture profile, which can be generated through vapor sorption (see Fig. 5; we refer to Fig. A.1 b&c and appendix: A.2 for a detailed description of the numerical model and the results).
Conclusions
We confirmed the existence and quantified the thickness of the micrometer-thick self-lubrication layer that forms via particle-fluid phase segregation during the high pressure extrusion of pastes. We show that this lubrication layer directly affects the tribological properties of the paste flow. We used realistic, plant-based materials inside an industrial-scale pellet extruder to evidence how the fundamental aspects of the lubrication layer interlinks industrial process variables. We validated our measurements through numerical simulations of liquid transport around non-uniformly wetted objects, and in doing so, unravel the fundamental mechanism behind a decades-old observation of the lubricating effect of water addition during pellet production [40].
@@ Conclusions
Our work reveals that the friction of an extrudate within a die is directly influenced by the thickness of the lubrication layer, providing insights for future monitoring of product quality, shape, and performance in food, feed, and pharmaceutical extrusion processes and adjacent extrusion disciplines which process inorganic ingredient streams like ceramics. This work introduces a framework to understand and utilize the role of extrusion parameters such as die geometry, particle size distribution, and ingredient composition. Furthermore, continuous monitoring of the phase segregation process during extrusion presents a novel strategy for process control, with immediate global industrial relevance. For instance, improved lubrication control can yield significant energy savings and enable more sustainable, high-performance extrusion. Real-time tracking of phase segregation may also enhance our understanding of the complex particle-liquid interactions, which we can subsequently use in quality control across a wide range of extrusion-based manufacturing processes.
@@ Acknowledgements.
We gratefully acknowledge discussions with the “Pelleting in Circular Agriculture” consortium members. We thank the Wageningen Electron Microscopy Center (WEMC) of Wageningen University & Research for access to facilities. We thank the Van der Waals-Zeeman Institute for providing access to the Keyence VK-X1000. We thank Zohreh Farmani and Jordi Rijpert for their help during the experimental investigation. We thank Remco Fokkink for helping with the development of software to read out the relevant sensors during the drying experiments. The authors thank Synchrotron SOLEIL for the ANATOMIX beamline time (proposal n° 20220272). ANATOMIX is an Equipment of Excellence (EQUIPEX) funded by the Investments for the Future program of the French National Research Agency (ANR), project NanoimagesX, grant no. ANR-11-EQPX-0031. This study was financially supported by The VICTAM Foundation, Agrifirm NWE B.V., DSM, Elanco Animal Health, Pelleting Technology Netherlands, Phileo S.I. Lesaffre, Topsector Agri & Food, Zetadec, and Wageningen University & Research, as partners in the project “Pelleting in Circular Agriculture” (project number: LWV1965)
@@ Appendix A Extended Data
Figure A.1: (a) A schematic representation illustrates the initial radial moisture distribution used as the initial condition in numerical modeling of the drying process for freshly produced pellets. The moisture ratio (unitless) represents the normalized mass decay over 900900900 seconds of drying (see Eq. 6). Numerical simulations provide a qualitative comparison of the effect of radial phase segregation on drying behavior, shown in both numerical simulations (b) and experimental drying data (c) for fresh (blue lines) and rehydrated pellets (yellow lines) using the moisture distributions as the initial condition (a). The difference in drying behavior is caused by radial phase segregation, specifically the formation of a lubrication layer during pellet extrusion.
We examined the pellet drying behavior, internal and external pellet microstructure, and surface topology of pellets produced under systematically varied process conditions, adjusting only the amount of added steam to the ingredient mixture when formulating the paste, thereby changing its initial water content prior to compaction, $\Delta\mathrm{mc}$. Our objective was to determine whether the self-lubricating liquid surface layer emerges during pellet extrusion. In the processing of pastes and other dense particulate suspensions, such phase segregation processes are also referred to as “spatial demixing,” “liquid migration,” “phase redistribution,” “drainage,” or “self-filtration” [12, 36]. According to our hypothesis, a thin lubricating film forms between the pellet and the metal die due to phase segregation (Fig. 2a), which fills the asperities between the two sliding surfaces and reducing the frictional contact area. As a result of the radial liquid migration, the moisture distribution across the pellets becomes non-uniform, with the exterior surface of the pellet facing the die wall experiencing increased moisture levels during extrusion compared to the pellet’s interior volume [25]. Through the combined circumstantial evidence presented below, we conclude that such phase migration occurs during pellet extrusion.
@@ A.1 Pellet production
A.2 Water distribution inside pellets: $\phi_{\rm liquid}$
In our numerical simulations, the initial moisture distribution $\phi_{\rm liquid}$ along the radial coordinate $r$ was described by a half-Gaussian function $M(r)$—a Gaussian centered at the pellet-air interface ($r_{p}$).
$\phi_{\rm liquid}\equiv M(r)=M_{\rm surf}\exp\left[-\frac{1}{2}\left(\frac{r-r_{p}}{\sigma}\right)^{2}\right]$ (2)
@@ A.2 Water distribution inside pellets: $\phi_{\rm liquid}$
Consequently, for the numerical model, mass transfer coefficient ($k_{c}$) and effective diffusion coefficient ($D_{\rm eff}$) were selected in a way that the transport rate at which water moves within the pellet (diffusion of water) is much lower than the rate of evaporation (mass transfer rate representing the change of liquid water on the surface into water vapor within the passing air). This is associated with a mass Biot number ($\mathrm{Bi_{m}}$) much larger than 111 (see Appendix C.1). With a mass Biot number much larger than 111, we ensure that the rate of evaporation at the surface does not limit the drying process, as the maximum evaporation rate is lower than the vapor capacity of the passing fresh air, as observed in the experimental data (see Fig. 2b).
@@
The qualitative comparison between the numerical model and the experimental drying data supports the hypothesis that a liquid surface layer emerges during pellets extrusion. Combining the quick initial drying of the fresh pellets together with the observation of air saturation (Fig. 2 b), we conclude that water must be present in close proximity to the pellet exterior surface as a result of liquid surface migration during pellet extrusion.
@@ A.3 Pellet microstructure
During pellet extrusion, the finely ground organic ingredients are bound into a cylindrical shaped pellet. Such organic ingredient streams deform easily at the typical process pressures, resulting in densification of the ingredient mixture, expelling pore fluid and enhancing the agglomeration process through various bond types [21]. The mechanical properties [61, 62], as well as the glass and melt transition temperatures of such organic plant derived products are strongly influenced by the local moisture content [6, 39]. Consequently, the local increase in moisture content due to the presence of a water rich lubrication layer at the pellets surface is expected to affect the surface microstructure of the pellet. Hence, we employed x-ray tomography at the synchrotron SOLEIL, utilizing the ANATOMIX beamline [63], to analyze the pellet microstructure. This analysis focused on identifying density gradients and structural changes associated with shear deformations and moisture gradients. Such gradients and structural changes can offer insights into the flow profile inside the extrusion die and the local moisture content during extrusion.
@@ A.3.1 A uniform radial pellet density profile
Figure A.2: Radial density profiles of pellets with a diameter of 6 mm, obtained through analysis of x-ray tomograms obtained at the ANATOMIX beamline, for two process conditions with (a) $\Delta\mathrm{mc}$ 0.0090.0090.009 (low water content) and (b) 0.0570.0570.057 (high water content) respectively. The light gray lines in (a) and (b) correspond to individual slices along the 3D volume. The black lines in (a) and (b) correspond to the average density profile over each of the slices within the stack (see Materials & Methods). The region marked by the gray area, corresponds to the area in which we observe a phase-contrast enhancement artifact. In this region the apparent density could not be determined accurately.
@@ A.3.3 Internal pellet deformation
It has been previously proposed that the ingredient stream moves as a plug inside the die channels [64], with a uniform flow velocity through the pellet die. A lubrication layer should facilitate such plug-flow and therefore, signs of internal deformations caused by velocity gradients should be absent. A qualitative analysis of x-ray tomograms confirms this absence of internal deformations along the longitudinal axis as indications of the flow direction (Fig. A.5). Additionally, the overall porosity is low (Fig. A.5), suggesting that the ingredients are forced into place, approaching unit density during pelletization. Indeed, the estimated process pressures during pelletization [65], are of similar order, or larger than the typical compressive strength of the particles [66] and hence the particles are compressed into a tightly packed pellet with a low porosity. These observations support the hypothesis of liquid surface migration, wherein water in between the particles, is expelled from the ingredient matrix into the pellet-die gap during pellet extrusion, leading to the formation of a lubrication layer that enables plug flow inside the pellet die.
Figure A.5: Longitudinal cross sections (x-ray tomograms obtained at the ANATOMIX beamline) of four pellets produced during the experimental trial at two different water contents after steam treatment, (a) two at low $\Delta\mathrm{mc}$ and (b) two at high $\Delta\mathrm{mc}$, 0.009 and 0.057 kg water per kg ingredients respectively. No indication for the flow direction is visible within the cross sections. The pellets flow as plugs through the pelleting die [64]. Scale bar indicates 0.50.50.5 cm.